Tardis offers a 3% discount to customers who pay cash at the time of sale and a 2% discount to customers who pay within the first 10 days of the month after sale. Past experience shows that cash collections from customers tend to occur in the following pattern: Cash collected at time of sale 55% Collected within cash discount period in first 10 days of month after sale 15% Collected after cash discount period in first month after month of sale 10% Collected after cash discount period in second month after month of sale 15% Never collected 5% Compute the total cash budgeted to be collected in March if sales forecasts are $370,000 for January, $420,000 for February, and $460,000 for March.
Answer #1
March
Cash collected from March Sales net of 3% discount $       2,45,410 =460000*55%*(100%-3%)
Cash collected from March Sales net of 2% discount $          67,620 =460000*15%*(100%-2%)
Cash collected from February Sales without discount $          42,000 =420000*10%
Cash collected from January Sales without discount $          55,500 =370000*15%
Total Cash Collections $      4,10,530